OpenStudy (anonymous):

Careful measurements have been made of Olympic sprinters in the 100-meter dash. A simple but reasonably accurate model is that a sprinter accelerates at 3.4 m/s2 for 3.30 s, then runs at constant velocity to the finish line.

OpenStudy (aum):

In problem makes two statements but does not ask any questions. But ... Sprinter starts from rest and accelerate at 3.4 m/s^2 for 3.30 s. u = 0 a = 3.4 m/s^2 t = 3.3 s v = ? v = u + at = 0 + 3.4 * 3.3 = 11.22 m/s So the sprinter accelerates to the speed of 11.22 m/s after which the sprinter runs at constant velocity to the finish line. Not knowing the question hopefully the answer is contained above.